Given a triangle ABC with an altitde AD and sides AB and AC measuring 5 units, prove that AD bisects BC.
Answer by RAY100(1637)

Triangle,,,ABC,,altitude AD, AB=AC=5,,,,,all GIVEN
.
AD perpendicular to BC,,,,,,,,,,,,,,,,Def of Altitude
.
ABD,,,and ACD are right triangles,,,,,,,,,def of rt triangle
.
.
AD =AD,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,Same Line
.
AB^2 -AD^2 = BD^2,,,,,,,,,,,,,,,,,,Pythagorean Theorem
.
AC^2 -AD^2 = CD^2,,,,,,,,,,,,,,,,,,Pythagorean Theorem
.
BD^2 =CD^2,,,,,,,,,,,,,,,,,,,,,,,,,arithmetic and substitution(AB+AC)
.
BD = CD,,,,,,,,,,,,,,,,,,,,,,,,,,,,square root
.
Therefore AD, bisects BC,,,,,,,,def of bisector
.